# Critical re-examination of a claimed challenge of Bohmian mechanics

### v3

@@ II Detailed analysis of the one-dimensional problem
Before moving to the full bidimensional Schrödinger equation (5), it might be useful here to remember the main outcomes of the simpler one-dimensional case of a step potential, in transient conditions, along the description of, e.g., [5, 6]. Consider the one-dimensional Schrödinger equation in a step potential:
@@
where $V_{0}$ is a positive constant and $\theta(x)$ the Heaviside function (1 if $x>0$ and 0 otherwise). Define as usual $k=\frac{\sqrt{2mE}}{\hbar}$ and $\kappa=\frac{\sqrt{2m(V_{0}-E)}}{\hbar}$. Then, take a rectangular wave packet, with $E<V_{0}$, at the initial time $t=0$: $\psi(x,0)=e^{ik_{0}x}\left(\theta(x+L)-\theta(x)\right)$, with $L\gg 1/k_{0}$. This wave-packet is not strictly monochromatic at $k_{0}$ because of the cutting edge at $x=0$ (the one at x=Lx=-L is not actually necessary in our case, as we are interested only in the transient phenomena at the raising edge: it will be eliminated below by letting $L\rightarrow\infty$). The packet travels towards the positive $x$ direction and reaches the step potential, at $x=0$, in the time interval $t\in[0,T]$, with $T=\frac{L}{v_{g}}$ (and $v_{g}$ the group velocity).
@@
with $z(x,t)\!=\!-\frac{k_{0}}{\kappa_{0}}\sqrt{\frac{m}{2\hbar t}}\left(x-iv_{0}t\right)e^{i\pi/4}$ and $v_{0}=\frac{\hbar\kappa_{0}}{m}$. We remind that $\frac{1}{2}{\rm erfc}\left(z(x,t)\right)\xrightarrow[t\rightarrow+\infty]{}1$, so that we recover in this limit the usual stationary solution $\psi_{\rm st}(x,t)=T(k_{0})e^{-\kappa_{0}x}e^{-iE_{0}t/\hbar}$, characterized by a null spatial-dependent phase ($\partial_{x}S=0$) and no current. Yet, this is not true during the transient period and in particular for $t\rightarrow 0^{+}$, when $z(x,t)\rightarrow-\sqrt{\frac{m}{2\hbar t}}\frac{k_{0}}{\kappa_{0}}xe^{i\pi/4}$.
If we keep only the leading term of the asymptotic expansion of erfc(z): ${\rm erfc}(z)\xrightarrow[|z|\rightarrow+\infty]{}\frac{e^{-z^{2}}}{\sqrt{\pi}z}+o(|z|^{-1})$, we obtain the time-dependent probability density (remark that $z^{2}$ is purely imaginary and equal to 1 in the $t=0^{+}$ limit):
$|\psi(x,t=0^{+})|^{2}\simeq\frac{1}{\pi}|T(k_{0})|^{2}e^{-2\kappa_{0}x}\frac{\kappa_{0}}{k_{0}}\frac{\hbar t}{2mx^{2}}$ (3)
Deriving Eq. (3) with respect to $t$ and integrating the result with respect to $x$, the continuity equation leads to the following expression for the current:
$j(x,t=0^{+})\simeq\frac{1}{\pi}|T(k_{0})|^{2}e^{-2\kappa_{0}x}\frac{\kappa_{0}}{k_{0}}\frac{\hbar}{2mx}$ (4)

### v4

@@ Appendix B The transient regime in the one-dimensional case
It might be useful here to remind the main outcomes of the one-dimensional case of a step potential, in transient conditions, along the description of, e.g., [4, 5]. Consider the one-dimensional Schrödinger equation in a step potential:
@@
where $V_{0}$ is a positive constant and $\theta(x)$ the Heaviside function (1 if $x>0$ and 0 otherwise). Define as usual $k=\frac{\sqrt{2mE}}{\hbar}$ and $\kappa=\frac{\sqrt{2m(V_{0}-E)}}{\hbar}$. Then, take a rectangular wave packet, with $E<V_{0}$, at the initial time $t=0$: $\psi(x,0)=e^{ik_{0}x}\left(\theta(x+L)-\theta(x)\right)$, with $L\gg 1/k_{0}$. This wave-packet is not strictly monochromatic at $k_{0}$ because of the cutting edge at $x=0$ (the one at x=Lx=-L is not actually necessary in our case, as we are interested only in the transient phenomena at the raising edge: it will be eliminated below by letting $L\rightarrow\infty$). The packet travels towards the positive $x$ direction and interacts with the step potential, at $x=0$, for $t\geq 0$.